OpenStudy (anonymous):

Transform the formula for the given variable: d = rt Solve for t. A. t = dr B. t = d + r C. t = d/r D. t = r/d

jimthompson5910 (jim_thompson5910):

Take d = rt and divide both sides by 'r' to isolate t d = rt d/r = rt/r d/r = t t = d/r So the answer is choice C

OpenStudy (anonymous):

Thanks (:

jimthompson5910 (jim_thompson5910):

you're welcome